# Field Matters: A lightweight LLM-enhanced Method for CTR Prediction

### v1

@@ Abstract
Click-through rate (CTR) prediction is a fundamental task in modern recommender systems. In recent years, the integration of large language models (LLMs) has been shown to effectively enhance the performance of traditional CTR methods. However, existing LLM-enhanced methods often require extensive processing of detailed textual descriptions for large-scale instances or user/item entities, leading to substantial computational overhead. To address this challenge, this work introduces LLaCTR, a novel and lightweight LLM-enhanced CTR method that employs a field-level enhancement paradigm. Specifically, LLaCTR first utilizes LLMs to distill crucial and lightweight semantic knowledge from small-scale feature fields through self-supervised field-feature fine-tuning. Subsequently, it leverages this field-level semantic knowledge to enhance both feature representation and feature interactions. In our experiments, we integrate LLaCTR with six representative CTR models across four datasets, demonstrating its superior performance in terms of both effectiveness and efficiency compared to existing LLM-enhanced methods. Our code is available at https://anonymous.4open.science/r/LLaCTR-EC46.
1 Introduction
@@
Click-through rate (CTR) prediction [1] is a core task in modern recommender systems (RS), with the objective of estimating the likelihood that a user will click on a specific item by modeling complex feature interactions. With the advent of Large Language Models (LLMs) and their remarkable abilities in content comprehension and semantic reasoning [2, 3], there has been a surge of interest in leveraging LLMs for CTR prediction. Recent studies mainly explored two paradigms: 1) LLMs as CTR predictors [4, 5, 6], where LLMs are either prompted or fine-tuned to directly perform CTR prediction; 2) LLM-enhanced CTR models [7, 8, 9, 10], where LLMs are leveraged to augment traditional CTR models by injecting semantic knowledge.
@@
Despite their encouraging performance, these approaches encounter significant practical limitations, particularly regarding computational efficiency and economic feasibility. The high computational demands of LLMs make their direct deployment for online CTR prediction largely impractical, as they introduce considerable inference latency that violates real-time serving requirements. Although the second paradigm alleviates inference latency by retaining conventional CTR models for online serving, it still suffers prohibitive training costs. Specifically, these methods typically operate at the instance level [7, 9, 11] or the user/item level [8, 12, 10], requiring LLMs to process detailed textual descriptions for large-scale data instances or user/item entities. Empirically, we find that existing LLM-enhanced CTR methods (e.g., KAR[8], LLM-CF[9], CTRL[7] and EASE [10]) require over 290 times (on average) more computation time than baseline models on typical Amazon Video Games and MovieLens-1M datasets with millions of interactions (Figure 4 and Appendix B.1). This computational burden becomes even more pronounced when scaling to industrial settings involving billions of instances. These challenges naturally motivate a critical research question: How can we leverage LLMs to enhance CTR models in a more efficient and cost-effective manner?
@@
To address this challenge, we explore a novel field-level enhancement paradigm that utilizes LLMs to extract semantic knowledge from feature fields to enhance CTR models. As illustrated in Figure 1, fields denote categories of features in CTR prediction. This paradigm offers two principal advantages: 1) Computational Efficiency: The number of fields is orders of magnitude lower than that of instances or entities (e.g., thousands vs. billions in industrial RS), yielding a drastic reduction in LLM computational overhead. 2) Crucial Semantic Information: Conventional CTR methods primarily utilize field IDs as categorical indicators, often neglecting the rich semantic information inherent in field descriptions. In fact, field semantics can significantly enhance both feature representation and feature interaction modeling — the two core components contributing the success of CTR prediction. For example, recognizing that the feature “4.7” belongs to the field “average rating” provides insight into item quality; while understanding the relationship between fields such as “user income” and “item price” can inform the importance of feature interactions. Although field-level semantic knowledge may not be as rich as cumbersome instance-level semantic information, its conciseness and importance offer substantial potential for improving model performance.
@@
Motivated by these analyses, we introduce LLaCTR, a novel Lightweight LLM-enhanced CTR method through field-level enhancement. LLaCTR comprises two key components:
@@
Figure 1: Current LLM-enhanced CTR paradigm versus our Field-level enhancement paradigm.
2 Preliminaries
Task Formulation. This work focuses on click-through rate (CTR) prediction [13, 14], a core task in many personalized services (e.g., recommender systems). Let $\mathcal{D}=\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{N}$ denote the training dataset of user-item historical interaction records, where $\mathbf{X}_{i}=[\mathbf{x}_{i1},\mathbf{x}_{i2},\ldots,\mathbf{x}_{iK}]$ represents the input features across $K$-field for the $i$-th instance, and $\mathbf{x}_{ik}$ denotes the feature of the $k$-th field for the instance. In general, $\mathbf{x}_{ik}$ is a vector that can represent categorical, multi-valued, numerical features, etc. The label $y_{i}\in\{1,0\}$ indicates whether the user has clicked on the item in this instance. The goal of CTR is to learn a model from $\mathcal{D}$ that can accurately predict the click probability of a new instance.
Traditional CTR Models. Traditional CTR models [15, 16, 17, 18, 19, 20] primarily adopt deep learning paradigms and can generally be abstracted into three core components:
@@ 2 Preliminaries
$\mathbf{E}_{i}=\text{EmbeddingLayer}(\mathbf{X}_{i}),$ (1)
where $\mathbf{E}_{i}=[\mathbf{e}_{i1},\mathbf{e}_{i2},\ldots,\mathbf{e}_{iK}]\in\mathbb{R}^{K\times D}$ represents the learned feature embeddings of the instance $i$, with $\mathbf{e}_{ik}$ denoting the D-dimensional embedding vector for $k$-th field feature.
@@
$\Phi(\mathbf{X}_{i})=\text{FeatureInteraction}(\mathbf{X}_{i},\mathbf{E}_{i}),$ (2)
where $\Phi(\cdot)$ is the hidden representation learned from feature interactions. Among various architectures, the most representative one is the factorization machine (FM) [15], which explicitly model bi-level feature interactions:
$\Phi_{FM}(\mathbf{X}_{i})={w_{0}}+\sum\limits_{k=1}^{K}{{\mathbf{w}_{k}}^{T}{\mathbf{x}_{ik}}}+\sum\limits_{k=1}^{K}{\sum\limits_{l=k+1}^{K}\varsigma({{{\mathbf{x}_{ik}{\mathbf{x}_{il}}^{T}\text{<}\textbf{e}_{ik},{\textbf{e}_{il}}\text{>}}}})},$ (3)
where $w_{0},\mathbf{w}_{k}$ denotes the learnable weights, and <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > denotes the inner product of two vectors; $\varsigma(.)$ denotes the sum of all elements in the matrix. For convenience, here we adopt vector notation to represent the formula of Factorization Machines (FM), which is mathematically equivalent to the original formulation presented in [15]. The features are explicitly bi-level interacted with $\mathbf{x}_{ik}{\mathbf{x}_{il}}^{T}$ and its contribution is controlled by their embeddings $\text{<}\mathbf{e}_{ik},\mathbf{e}_{il}\text{>}$.
Building on FM, recent works have explored various extensions [21, 22, 23], such as introducing field-aware matrices [24, 17, 18] or incorporating neural layers [16, 25]. Meanwhile, other studies have proposed novel architectures, such as convolution operator [26, 27] and self-attention [28, 29], to better capture complex feature interactions.
@@
$\hat{y_{i}}=\sigma[\text{PredictionLayer}(\Phi(\mathbf{X}_{i}))],$ (4)
@@
where $\sigma(\cdot)$ is the sigmoid function. The model is typically optimized with the binary cross-entropy (BCE) loss [30] with:
$\mathcal{L}_{BCE}(y,\hat{y})=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log\left(\hat{y_{i}}\right)+(1-y_{i})\log\left(1-\hat{y_{i}}\right)\right],$ (5)
@@
While these traditional CTR models have made significant progress in the past few decades, they are often plagued by the issue of semantic information loss. Specifically, Some methods have utilized field IDs as categorical indicators [24, 17, 18], but the rich semantic information inherent in field descriptions has not been exploited. This limitation motivates us to leverage LLMs to enrich the traditional CTR models with field-level knowledge.
@@
LLM-enhanced CTR Models. To fully exploit semantic knowledge, Large Language Models (LLMs) [2, 3] have been extensively investigated for enhancing traditional CTR models [8, 7, 9, 10, 31]. As illustrated in Figure 1, existing LLM-enhanced CTR methods primarily operate at the instance level or user/item level. Typically, these approaches first organize the features of users, items, or instances into textual descriptions, which are then provided as prompts to the LLM for reasoning or summarization. The resulting semantic knowledge, often at the user/item or instance level, is further encoded into semantic embeddings to augment traditional CTR models (see Section 5 for details). However, due to the large scale of users/items and instances in real-world applications, such strategies incur prohibitive computational costs during training, inference and encoding. For example, on representative datasets such as Amazon Video Games and MovieLens-1M, the total training time of existing LLM-enhanced CTR methods (e.g., KAR [8], LLM-CF [9], CTRL [7] and EASE [10]) is over 290 times (on average) that of standard CTR models on average (see Appendix B.1 for details). Moreover, the complex and voluminous semantic knowledge generated at the user/item or instance level is often difficult for traditional CTR models to effectively assimilate and utilize. Given these limitations, it is imperative to explore new paradigms for integrating semantic knowledge into CTR models.
3 Methodology
Figure 2: The overall framework of proposed LLaCTR. It contains 1) SSFT: leveraging self-supervised field-feature fine-tuning to improve LLMs’ ability to capture field semantics; 2) FRE: leveraging field semantic embeddings to guide the learning of the feature representation; 3) FIE: leveraging field semantic embeddings to enhance the models of feature interactions.
@@ 3 Methodology
In this section, we introduce LLaCTR, which enhances CTR prediction through field-level semantic knowledge from LLMs. As shown in Figure 2, LLaCTR first employs self-supervised field-feature fine-tuning to improve LLMs’ ability to capture field semantics (Section 3.1), and then leverage the field knowledge distilled from LLMs to enhance both feature representation (Section 3.2) and feature interactions (Section 3.3).
3.1 Self-supervised Field-feature Fine-tuning (SSFT)
@@
LLMs are pre-trained on open-domain natural language corpora [2, 3], which often lack domain-specific knowledge relevant to CTR prediction. This limitation can hinder their ability to distill high-quality field semantic knowledge. Specifically, LLMs may not fully understand the meaning of field descriptions in CTR tasks or their relationships with features. To address this, we propose a self-supervised fine-tuning strategy that prompts LLMs to predict the field to which a given feature belongs, thereby enhancing their understanding of field semantic.
@@
$\mathcal{L}_{LG}(\mathcal{P}(f^{k}_{j}),F_{k};\theta)=-\log P_{LLM}\left(F_{k}\mid\mathcal{P}(f^{k}_{j})\right),$ (6)
@@
Besides the language generative loss, we find the contrastive loss [32] is also effective:
$\mathcal{L}_{CL}(\mathcal{P}(f^{k}_{j}),F_{k};\theta)=-\log\frac{\exp(cos(En(\mathcal{P}(f_{j}^{k})),En(F_{k}))/\tau)}{\sum_{l=1}^{K}\exp(cos(En(\mathcal{P}(f_{j}^{k})),En(F_{l}))/\tau)},$ (7)
where $cos(\cdot)$ denotes the cosine similarity function, $\tau$ is the temperature parameter and $En(\cdot)$ are the LLM’s encodings of the language descriptions, respectively.
The contrastive loss is designed to align the semantic embeddings of prompts with their corresponding answers, thereby effectively injecting feature-field correlation knowledge in embeddings. Since the contrastive loss operates directly on embeddings, which are subsequently used to enhance CTR models, this work prioritizes the use of contrastive loss and empirically find slightly better performance than generative loss (Appendix B.2).
After fine-tuning, the LLM’s encodings of field descriptions are distilled:
$\textbf{h}_{k}=En(F_{k}).$ (8)
@@
Figure 3: The template of self-supervised tuning.
Notably, such field-level operations are highly efficient, as the LLMs require to process only a small number of fields (typically less than 100) rather than millions of instances or user/item pairs in the experimental datasets. This difference in scale is even more pronounced in real-world industrial recommender systems, where there may be thousands of fields, but hundreds of millions of users/items and billions of instances. Some readers may be concerned about the potential inefficiency of self-supervised fine-tuning. However, this process is also lightweight, as it only requires sampling a small number of features per field for training. For example, sampling just 500 features per field is sufficient to achieve significantly better performance compared to recent LLM-enhanced methods, while incurring substantially lower training time. Empirical comparisons of efficiency can be found in Figure 4. The detailed time complexity analyses of LLaCTR and other LLM-enhanced baselines are provided in Appendix B.1.
3.2 Feature Representation Enhancement (FRE)
The quality of feature representations is critical for the success of CTR models. However, existing methods often overlook the semantic knowledge contained in field descriptions. To address this, we transfer the field semantic knowledge distilled from LLMs to enhance feature embeddings.

### v2

@@ Abstract.
Click-through rate (CTR) prediction is a fundamental task in modern recommender systems. In recent years, the integration of large language models (LLMs) has been shown to effectively enhance the performance of traditional CTR methods. However, existing LLM-enhanced methods often require extensive processing of detailed textual descriptions for large-scale instances or user/item entities, leading to substantial computational overhead. To address this challenge, this work introduces LLaCTR, a novel and lightweight LLM-enhanced CTR method that employs a field-level enhancement paradigm. Specifically, LLaCTR first utilizes LLMs to distill crucial and lightweight semantic knowledge from small-scale feature fields through self-supervised field-feature fine-tuning. Subsequently, it leverages this field-level semantic knowledge to enhance both feature representation and feature interactions. In our experiments, we integrate LLaCTR with six representative CTR models across four datasets, demonstrating its superior performance in terms of both effectiveness and efficiency compared to existing LLM-enhanced methods. Our code is available at https://github.com/istarryn/LLaCTR.
Large language Model, CTR Prediction, Recommender Systems
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792387isbn: 979-8-4007-2307-0/2026/04submissionid: rfp1942ccs: Information systems Recommender systems
1. Introduction
@@
Click-through rate (CTR) prediction (Yang and Zhai, 2022) is a core task in modern recommender systems (RS) (Chen et al., 2023; Zhang et al., 2025), with the objective of estimating the likelihood that a user will click on a specific item by modeling complex feature interactions. With the advent of Large Language Models (LLMs) and their remarkable abilities in content comprehension and semantic reasoning (Achiam et al., 2023; Dubey et al., 2024), there has been a surge of interest in leveraging LLMs for CTR prediction. Recent studies mainly explored two paradigms: 1) LLMs as CTR predictors (Geng et al., 2022; Bao et al., 2023b; Geng et al., 2024), where LLMs are either prompted or fine-tuned to directly perform CTR prediction; 2) LLM-enhanced CTR models (Li et al., 2023; Xi et al., 2024; Sun et al., 2024; Qiu et al., 2024), where LLMs are leveraged to augment traditional CTR models by injecting semantic knowledge.
@@
Despite their encouraging performance, these approaches encounter significant practical limitations, particularly regarding computational efficiency and economic feasibility. The high computational demands of LLMs make their direct deployment for online CTR prediction largely impractical, as they introduce considerable inference latency that violates real-time serving requirements. Although the second paradigm alleviates inference latency by retaining conventional CTR models for online serving, it still suffers prohibitive training costs. Specifically, these methods typically operate at the instance level (Li et al., 2023; Sun et al., 2024; Lin et al., 2024a) or the user/item level (Xi et al., 2024; Wang et al., 2024; Qiu et al., 2024), requiring LLMs to process detailed textual descriptions for large-scale data instances or user/item entities. Empirically, we find that existing LLM-enhanced CTR methods (e.g., KAR(Xi et al., 2024), LLM-CF(Sun et al., 2024), CTRL(Li et al., 2023) and EASE (Qiu et al., 2024)) require over 290 times (on average) more computation time than baseline models on typical Amazon Video Games and MovieLens-1M datasets with millions of interactions (cf. Figure 1). These challenges naturally motivate a critical research question: How can we leverage LLMs to enhance CTR models in a more efficient and cost-effective manner?
Figure 1. Empirical efficiency study on representative conventional CTR models and LLM-enhanced CTR models. The multiple increase in time cost is reported based on WuKong.
@@
To address this challenge, we explore a novel field-level enhancement paradigm that utilizes LLMs to extract semantic knowledge from feature fields to enhance CTR models. As illustrated in Figure 2, fields denote categories of features in CTR prediction. This paradigm offers two principal advantages: 1) Computational Efficiency: The number of fields is orders of magnitude lower than that of instances or entities (e.g., hundreds vs. millions in the real RS), yielding a drastic reduction in LLM computational overhead. 2) Crucial Semantic Information: Conventional CTR methods primarily utilize field IDs as categorical indicators, often neglecting the rich semantic information inherent in field descriptions. In fact, field semantics can significantly enhance both feature representation and feature interaction modeling — the two core components contributing the success of CTR prediction. For example, recognizing that the feature “4.7” belongs to the field “average rating” provides insight into item quality; understanding the relationship between fields such as “user income” and “item price” can inform the importance of feature interactions. Although field-level semantic knowledge may not be as rich as cumbersome instance-level semantic information, its conciseness and importance offer substantial potential for improving model performance.
@@
Motivated by these analyses, we introduce LLaCTR, a novel Lightweight LLM-enhanced CTR method through field-level enhancement, comprising two key components:
@@
Overall, this work makes the following contributions:
  • We identify the limitations of existing LLM-based CTR methods and advocate for leveraging field-level semantic knowledge to efficiently enhance traditional CTR models.
  • We propose a novel LLM-enhanced CTR method, LLaCTR, which employs self-supervised fine-tuning to distill high-quality field knowledge from LLMs and integrates this knowledge to improve both feature representation and feature interaction modeling.
  • We conduct extensive experiments to demonstrate that LLaCTR significantly improves the accuracy of traditional CTR models while being highly efficient and cost-effective.
Figure 2. Current LLM-enhanced CTR paradigm versus our field-level enhancement paradigm.
2. Preliminaries
2.1. Task Formulation
This work focuses on click-through rate (CTR) prediction (Lin et al., 2023; Wang et al., 2022), a core task in many personalized services (e.g., recommender systems). Let $\mathcal{D}=\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{N}$ denote the training dataset of user-item historical interaction records, where $\mathbf{X}_{i}=[\mathbf{x}_{i1},\mathbf{x}_{i2},\ldots,\mathbf{x}_{iK}]$ represents the input features across $K$-field for the $i$-th instance, and $\mathbf{x}_{ik}$ denotes the feature of the $k$-th field for the instance. In general, $\mathbf{x}_{ik}$ is a vector that can represent categorical, multi-valued, numerical features, etc. The label $y_{i}\in\{1,0\}$ indicates whether the user has clicked on the item in this instance. The goal of CTR is to learn a model from $\mathcal{D}$ that can accurately predict the click probability of a new instance.
2.2. Traditional CTR Models
Traditional CTR models (Rendle, 2010; Guo et al., 2017; Pan et al., 2018; Sun et al., 2021; Mao et al., 2023b; Zhang et al., 2024) primarily adopt deep learning paradigms and can generally be abstracted into three core components:
@@ 2.2. Traditional CTR Models
(1) $\mathbf{E}_{i}=\text{EmbeddingLayer}(\mathbf{X}_{i}),$
where $\mathbf{E}_{i}=[\mathbf{e}_{i1},\mathbf{e}_{i2},\ldots,\mathbf{e}_{iK}]\in\mathbb{R}^{K\times D}$ represents the learned feature embeddings of the instance $\mathbf{X}_{i}$, with $\mathbf{e}_{ik}$ denoting the D-dimensional embedding vector for $k$-th field feature.
@@
(2) $\Phi(\mathbf{X}_{i})=\text{FeatureInteraction}(\mathbf{X}_{i},\mathbf{E}_{i}),$
where $\Phi(\cdot)$ is the hidden representation learned from feature interactions. Among various architectures, the most representative one is the factorization machine (FM) (Rendle, 2010), which explicitly model bi-level feature interactions:
(3) $\small\Phi_{FM}(\mathbf{X}_{i})={w_{0}}+\sum\limits_{k=1}^{K}{{\mathbf{w}_{k}}^{T}{\mathbf{x}_{ik}}}+\sum\limits_{k=1}^{K}{\sum\limits_{l=k+1}^{K}\varsigma({{{\mathbf{x}_{ik}{\mathbf{x}_{il}}^{T}\text{<}\textbf{e}_{ik},{\textbf{e}_{il}}\text{>}}}})},$
where $w_{0},\mathbf{w}_{k}$ denotes the learnable weights, and $\cdot$$\cdot$> denotes the inner product of two vectors; $\varsigma(.)$ denotes the sum of all elements in the matrix. For convenience, here we adopt vector notation to represent the formula of Factorization Machines (FM), which is mathematically equivalent to the original formulation presented in (Rendle, 2010). The features are explicitly bi-level interacted with $\mathbf{x}_{ik}{\mathbf{x}_{il}}^{T}$ and its contribution is controlled by their embeddings $\text{<}\mathbf{e}_{ik},\mathbf{e}_{il}\text{>}$.
Building on FM, recent works have explored various extensions (Xiao et al., 2017; Lu et al., 2021; Tao et al., 2020), such as introducing field-aware matrices (Juan et al., 2016; Pan et al., 2018; Sun et al., 2021) or incorporating neural layers (Guo et al., 2017; Zhang et al., 2019). Meanwhile, other studies have proposed novel architectures, such as convolution operator (Liu et al., 2019a; Li et al., 2019) and self-attention (Zhou et al., 2018, 2019), to better capture complex feature interactions.
@@
(4) $\hat{y_{i}}=\sigma[\text{PredictionLayer}(\Phi(\mathbf{X}_{i}))],$
@@
where $\sigma(\cdot)$ is the sigmoid function. The model is typically optimized with the binary cross-entropy (BCE) loss (Mao et al., 2023a) with:
[…]
@@
While these traditional CTR models have made significant progress in the past few decades, they often suffer from the semantic information loss. Specifically, some methods rely solely on field IDs as categorical indicators (Juan et al., 2016; Pan et al., 2018; Sun et al., 2021), neglecting the rich semantic information in field descriptions. This limitation motivates our use of LLMs to enhance traditional CTR models with field-level semantic knowledge.
Figure 3. The overall framework of proposed LLaCTR.
2.3. LLM-enhanced CTR Models
@@
To fully exploit semantic knowledge, LLMs have been extensively investigated for enhancing traditional CTR models. As illustrated in Figure 2, existing LLM-enhanced CTR methods primarily operate at the instance level or user/item level. Typically, these approaches first organize the features of users, items, or instances into textual descriptions, which are then provided as prompts to the LLM for reasoning or summarization. The resulting semantic knowledge, often at the user/item or instance level, is further encoded into semantic embeddings to augment traditional CTR models.
However, due to the large scale of users/items and instances in real-world applications, such strategies incur prohibitive computational costs during training, inference and encoding. For example, on representative datasets such as Amazon Video Games and MovieLens-1M, the total training time of existing LLM-enhanced CTR methods (e.g., KAR (Xi et al., 2024), LLM-CF (Sun et al., 2024), CTRL (Li et al., 2023) and EASE (Qiu et al., 2024)) is over 290 times (on average) that of standard CTR models on average (as shown in Figure 1). Moreover, the complex and voluminous semantic knowledge generated at the user/item or instance level is often difficult for traditional CTR models to effectively assimilate and utilize. Given these limitations, it is imperative to explore new paradigms for integrating semantic knowledge into CTR models.
3. Methodology
@@
In this section, we introduce LLaCTR, which enhances CTR prediction through field-level semantic knowledge from LLMs. As shown in Figure 3, LLaCTR first employs self-supervised field-feature fine-tuning to improve LLMs’ ability to capture field semantics (SSFT), and then leverages the field knowledge distilled from LLMs to enhance both feature representation (FRE) and feature interactions (FIE).
3.1. Self-supervised Field-feature Fine-tuning (SSFT)
@@
LLMs are pre-trained on open-domain natural language corpora (Achiam et al., 2023; Dubey et al., 2024), which often lack domain-specific knowledge relevant to CTR prediction. This limitation can hinder their ability to distill high-quality field semantic knowledge. Specifically, LLMs may not fully understand the meaning of field descriptions in CTR tasks or their relationships with features. To address this, we propose a self-supervised fine-tuning strategy that prompts LLMs to predict the field to which a given feature belongs, thereby enhancing their understanding of field semantics.
@@
(6) $\mathcal{L}_{LG}(\mathcal{P}(f^{k}_{j}),F_{k};\theta)=-\log P_{LLM}\left(F_{k}\mid\mathcal{P}(f^{k}_{j})\right),$
@@
Besides the language generative loss, we find the contrastive loss (He et al., 2020) is also effective:
(7) $\small\mathcal{L}_{CL}(\mathcal{P}(f^{k}_{j}),F_{k};\theta)=-\log\frac{\exp(cos(\mathcal{E}(\mathcal{P}(f_{j}^{k})),\mathcal{E}(F_{k}))/\tau)}{\sum_{l}\exp(cos(\mathcal{E}(\mathcal{P}(f_{j}^{k})),\mathcal{E}(F_{l}))/\tau)},$
where $cos(\cdot)$ denotes the cosine similarity function, $\tau$ is the temperature parameter and $\mathcal{E}(\cdot)$ are the LLM’s encodings of the language descriptions, respectively.
The contrastive loss is designed to align the semantic embeddings of prompts with their corresponding answers, thereby effectively injecting feature-field correlation knowledge in embeddings. Since the contrastive loss operates directly on embeddings, which are subsequently used to enhance CTR models, this work prioritizes the use of contrastive loss and empirically find slightly better performance than generative loss (see Appendix B.2.1 for details).
After fine-tuning, the LLM’s encodings of field descriptions are extracted:
(8) $\textbf{h}_{k}=\mathcal{E}(F_{k}).$
@@
Figure 4. The template of self-supervised fine-tuning.
Notably, such field-level operations are highly efficient, as the LLMs require to process only a small number of fields (typically less than 100) rather than millions of instances or user/item pairs in the experimental datasets. Some readers may be concerned about the potential inefficiency of self-supervised fine-tuning. However, this process is also lightweight, as it only requires sampling a small number of features per field for training. For example, sampling just 500 features per field is sufficient to achieve significantly better performance compared to recent LLM-enhanced methods, while incurring substantially lower training time. Empirical comparisons of efficiency can be found in Figure 5. The detailed time efficiency analyses of LLaCTR and other LLM-enhanced baselines are provided in Section 4.2.2.
3.2. Feature Representation Enhancement (FRE)
The quality of feature representations is critical for the success of CTR models. However, existing methods often overlook the semantic knowledge contained in field descriptions (Lin et al., 2024a; Xi et al., 2024). To address this, we transfer the field semantic knowledge distilled from LLMs to enhance feature embeddings.